OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (anonymous):

y = 3x - 11

OpenStudy (anonymous):

here is the explanation since it is parallel to y = 3x -6 the gradient = 3 so the general equation will be y = 3x + c now to find c substitute the coordinates and you get -2 = 9 + c c = -11 therefore the equation is y = 3x -11

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

glad to be of assistance